Let (sin x) + (cos x) = 1.24 and csc x + sec x = 6.58. Which is (sin x)(cos x) ?

A. .13456
B. .16
C. .1884
D. .14
1
Expert's answer
2013-09-09T10:10:14-0400

sin x + cos x = 1.24
csc x + sec x = 6.58

1/cosx + 1/sin x = 6.58
(sin x +cos x)/(sinx*cos x) = 6.58

(sin x)*(cos x)= 1.24/6.58 = 0.1884

Answer: C